OpenStudy (anonymous):

simplify the expression 3735

TheSmartOne (thesmartone):

It looks simplified...

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

beating up a dead horse doesn't make it deader

OpenStudy (anonymous):

i know but that ont what the school is saying heres my choices 3^35 , 3^12, 139 or 9

OpenStudy (jdoe0001):

ahemm exponentialized then.... not sure if one could call that "simplified" though

OpenStudy (anonymous):

thats the question exactly

OpenStudy (anonymous):

Well 3^35 is 50031545098999707, 3^12 is 531441 and then there is 139 and 9. None of those equal 3735, so the answer is None of the Above...

OpenStudy (anonymous):

thats not one of my choices

OpenStudy (anonymous):

I know. It must be a bad question... Not much you can do about it

OpenStudy (anonymous):

i wonder if they meant maybe  this maybe it way a typo sometimes that happens would any of the choices that i wrote be correct if it is wrote out that way

OpenStudy (jdoe0001):

\(\bf 3735\implies 3\cdot 3\cdot 5\cdot 83({\color{brown}{ \textit{prime number}}})\) so once you get a prime there.... no dice on getting any exponentialized version of factors, let alone with base of "3"

OpenStudy (anonymous):

i know it doesnt make sense my school does this sometimes my aunt is teacher and she couldnt figure it out either